A capacitive time constant of an RC series circuit is 8.75 ms. The resistance is 2.50...

  1. A capacitive time constant of an RC series circuit is 8.75 ms. The resistance is 2.50 kΩ.
  1. Calculate the capacitance of the circuit
Answer #1

Time Constant of RC circuit

\tau= C ×R = 8.75×10-3 ,

Given R = 2.50 × 103 ohm

Hence, C = \tau /R = (8.75×10-3)/(2.50×103)

C = 3.5×10-6 F

or, C = 3.5 \mu F (Since, \mu = 10-6 or micro )